OpenStudy (anonymous):

Can someone please explain to me home to work this logarithmic equation, I have finals Monday!!!! Solve: 2 log3 (x+2)=log3 9+2

OpenStudy (anonymous):

typo *how

OpenStudy (anonymous):

deja vu, that strange feeling you get when you think you have been there before

OpenStudy (anonymous):

I'm sorry lol I just don't get it

OpenStudy (anonymous):

\[2\log_3(x+2)=\log_3(9)+2\] and if it says in the problem \[2\log_3(x+2)=\log_39+2\] it means the same thing first thing we do is change \[\log_39\] to 2 because \[3^2=2\] is that part ok?

OpenStudy (anonymous):

O k.....

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\] so \[\log_3(9)=2\iff 3^2=9\]

OpenStudy (anonymous):

okay got u

OpenStudy (anonymous):

so now we have \[2\log_3(x+2)=2+2=4\] divide both sides by 2 to get \[\log_3(x+2)=2\] so far so good?

OpenStudy (anonymous):

Where did the 2+2=4 come from?

OpenStudy (anonymous):

original problem had \[\log_39+2\] on the right, and since \[\log_39=2\] we get \[\log_39+2=2+2=4\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

now divide by 2 to get \[\log_3(x+2)=2\]

OpenStudy (anonymous):

and rewrite in "equivalent exponential form" meaning change \[\log_b(x)=y\text { to } b^y=x\] so turn this into \[3^2=x+2\]

OpenStudy (anonymous):

then solve for x by subtracting 2 from both sides

OpenStudy (anonymous):

Okay thank you much :)

OpenStudy (anonymous):

yw